# Stackelberg-Pareto Synthesis with Quantitative Reachability Objectives

### v2

@@ Abstract
In this paper, we deepen the study of two-player Stackelberg games played on graphs in which Player 00 announces a strategy and Player $1$, having several objectives, responds rationally by following plays providing him Pareto-optimal payoffs given the strategy of Player 00. The Stackelberg-Pareto synthesis problem, asking whether Player 00 can announce a strategy which satisfies his objective, whatever the rational response of Player $1$, has been recently investigated for $\omega$-regular objectives. We solve this problem for weighted graph games and quantitative reachability objectives such that Player 00 wants to reach his target set with a total cost less than some given upper bound. We show that it is NEXPTIME-complete, as for Boolean reachability objectives.
@@ 2.2 Stackelberg-Pareto Synthesis Problem
In an SP game, Player 0 wishes to minimize the value of a play with respect to the usual order $<$ on ${\rm Nature}$ extended to $\overline{{\rm Nature}}$ such that $n<\infty$ for all $n\in{\rm Nature}$. To compare the costs of Player $1$, the following component-wise order is introduced. Let $c,c^{\prime}\in\overline{{\rm Nature}}^{t}$ be two costs, we say that $c\leq c^{\prime}$ if $c_{i}\leq c_{i}^{\prime}$ for all $i\in\{1,\ldots,t\}$. Moreover, we write $c<c^{\prime}$ if $c\leq c^{\prime}$ and $c\neq c^{\prime}$. Notice that the order defined on costs is not total. Given two plays with respective costs $c$ and $c^{\prime}$, if $c<c^{\prime}$, then Player $1$ prefers the play with lower cost $c$.
@@
Figure 1: Arena $A$ (on the left) – Witness tree (on the right)
@@
Interestingly, the Boolean version of this game does not admit any solution. In this case, given a target, the player’s goal is simply to visit it (and not to minimize the cost to reach it). That is, the Boolean version is equivalent to the quantitative one with all weights and the bound $B$ put to zero. In the example, the play $v_{0}v_{1}v_{2}^{\omega}$ is Pareto-optimal (with visits to $T_{1}$ and $T_{3}$), whatever the strategy of Player 00, and this play does not visit Player 00’s target.
@@ 2.3 Tools
Let $X=\{e\in E\mid w(e)\geq 2\}$ be the set of edges of $A$ with weight at least 2. For each $e\in X$, we replace $e$ by a succession of $w(e)-1$ new vertices linked by $w(e)$ new edges of weight 1, and for each $e\in E\setminus X$, we keep $e$ unmodified. In this way, we get a directed graph $(V^{\prime},E^{\prime})$ with a weight function $w^{\prime}\colon E^{\prime}\rightarrow\{0,1\}$ and with a size $|V^{\prime}|\leq|V|\cdot W$. Notice that if $X=\emptyset$, the arena is already binary. The new arena $A^{\prime}=(V^{\prime},V^{\prime}_{0},V^{\prime}_{1},E^{\prime},v_{0},w^{\prime})$ has the same initial vertex $v_{0}$ as $A$, and a partition $V^{\prime}_{0}\cup V^{\prime}_{1}$ of $V^{\prime}$ such that each new vertex is added [^5] Each new vertex could be added to $V_{1}$ as it has a unique successor. to $V_{0}$, hence $V_{0}\subseteq V^{\prime}_{0}$, $V_{1}=V^{\prime}_{1}$. The new SP game $G^{\prime}=(A^{\prime},T_{0},\ldots,T_{t})$ keeps the same targets as in $G$.
@@ 3.2 Improving a Solution in a Subgame
  • If $h\not\rho$̸, then $\rho$ is also consistent with $\sigma_{0}^{\prime}$, thus there exists $c^{\prime}\in C_{\sigma_{0}^{\prime}}$ such that $c^{\prime}\leq c=\textsf{cost}({\rho})$ by definition of $C_{\sigma_{0}^{\prime}}$.
@@ 3.3 Eliminating a Cycle in a Witness
Then the strategy $\sigma_{0}^{\prime}=\sigma_{0}[\rho_{\leq m}\rightarrow\sigma_{0|\rho_{\leq n}}]$ is a solution in $\mbox{SPS}({G},{B})$ such that $\sigma_{0}^{\prime}\preceq\sigma_{0}$. Moreover, if $\sigma_{0}^{\prime}\simeq\sigma_{0}$, then $\sigma^{\prime}_{0}$ has a witness set $\textsf{Wit}_{\sigma^{\prime}_{0}}$ such that $\textsf{length}({\textsf{Wit}_{\sigma^{\prime}_{0}}})<\textsf{length}({\textsf{Wit}_{\sigma_{0}}})$.
Given a witness $\rho$, the first condition of the lemma means that $\rho_{[m,n]}$ is a cycle and that it appears before the last visit of a target by $\rho$. The second one says that $\rho_{\leq m}\sim\rho_{\leq n}$, i.e., no new target is visited and no branching point is crossed from history $\rho_{\leq m}$ to history $\rho_{\leq n}$. The third one says that if $\rho_{\leq m}$ does not visit Player 00’s target, then the cycle $\rho_{[m,n]}$ must have a null weight. The new strategy $\sigma^{\prime}_{0}$ is obtained from $\sigma_{0}$ by playing after $\rho_{\leq m}$ as playing after $\rho_{\leq n}$ (thus deleting the cycle $\rho_{[m,n]}$). See Figure 3.
$v_{0}$$\rho$$G$$\rho_{m}$$\rho_{n}$
@@
We first prove that $\sigma^{\prime}_{0}\preceq\sigma_{0}$. Let $c\in C_{\sigma_{0}}$ and $\pi\in\textsf{Wit}_{\sigma_{0}}$ be a witness with $\textsf{cost}({\pi})=c$. Let us prove that there exists $c^{\prime}\in C_{\sigma^{\prime}_{0}}$ such that $c^{\prime}\leq c$. If $g\not\pi$̸, then $\pi\in\textsf{Play}_{\sigma^{\prime}_{0}}$ by definition of $\sigma^{\prime}_{0}$, and thus there exists $c^{\prime}\in C_{\sigma^{\prime}_{0}}$ such that $c^{\prime}\leq c$ by definition of $C_{\sigma^{\prime}_{0}}$. If $g\pi$, as $g\sim h$ by hypothesis, we have that $\textsf{Wit}_{\sigma_{0}}(g)=\textsf{Wit}_{\sigma_{0}}(h)$ and no new target is visited from $g$ to $h$. It follows that $h\pi$ and $\bar{\pi}\in\textsf{Play}_{\sigma^{\prime}_{0}}$ with $\textsf{cost}({\bar{\pi}})=\textsf{cost}({\pi})-w(\rho_{[m,n]})\leq\textsf{cost}({\pi})$. By definition of $C_{\sigma^{\prime}_{0}}$, there exists $c^{\prime}\in C_{\sigma^{\prime}_{0}}$ such that $c^{\prime}\leq\textsf{cost}({\bar{\pi}})$, and therefore $c^{\prime}\leq\textsf{cost}({\pi})=c$. Hence $\sigma^{\prime}_{0}\preceq\sigma_{0}$.
@@
  • In the second case, we consider $\pi\in\textsf{Play}_{\sigma_{0}}$ such that $\bar{\pi}=\pi^{\prime}$. Notice that $\textsf{cost}({\pi})=c^{\prime}$ because $\rho_{[m,n]}$ has a null weight and $g\sim h$ (in particular, $\textsf{cost}({g})=\textsf{cost}({h})$, that is, no new Player 1’s target is visited from $g$ to $h$). We get that $c^{\prime}=\textsf{cost}({\pi})\in C_{\sigma_{0}}$ since $\sigma^{\prime}_{0}\preceq\sigma_{0}$. It follows that $\textsf{val}(\pi)\leq B$ as $\sigma_{0}\in\mbox{SPS}({G},{B})$. As $g\sim h$ (in particular, $\textsf{val}(g)=\textsf{val}(h)$), as $\textsf{val}(g)=\infty$, $\pi$ visits Player 00’s target after $h$, thus not along $\rho_{[m,n]}$. It follows that $\textsf{val}(\pi^{\prime})\leq B$. Therefore $\sigma^{\prime}_{0}\in\mbox{SPS}({G},{B})$.
@@ 4 Bounding Pareto-Optimal Payoffs
The theorem is proved by induction on the dimension $t$, with the calculation of a function $f({B},{t})$ depending on both $B$ and $t$, that bounds the components $c^{\prime}_{i}\neq\infty$. That is, in Theorem 4.1, Equation (2) is replaced by
@@ 4.1 Dimension One
Notice that $f({B},{1})$ satisfies (4) when $t=1$. Indeed, we have
@@ 4.2 Bounding the Pareto-Optimal Costs by Induction
The idea of the proof is as follows. If there exists $c\in C_{\sigma_{0}}$ such that for some $i\in\{1,\ldots,t+1\}$, $c_{i}$ does not satisfy (7), then we consider a witness $\rho$ with $\textsf{cost}({\rho})=c$ and the history $h$ of minimal length such that $h\rho$ and $w(h)=\max\{c_{min},B\}+1$. It follows that the subgame $G_{|h}$ has a smaller dimension and we can thus apply the induction hypothesis in $G_{|h}$ with $B=0$ as $h$ has already visited Player 00’s target. Hence, by (3), we get a better solution in the subgame $G_{|h}$, and then a better solution in the whole game $G$ by Lemma 3.1.
@@
[…]
@@
  • and with the targets visited by $\bar{h}$ removed from Player $1$’s set of targets, and if $k<t$ with $t-k$ additional targets that are copies of some remaining (non-empty) targets in a way to have exactly $t$ targets for Player $1$.
@@
We can thus apply the induction hypothesis: by Equation (3), there exists $\bar{\tau}_{0}\in\mbox{SPS}({\bar{G}},{0})$ without cycles such that $\bar{\tau}_{0}\preceq\bar{\sigma}_{0}$ and for all $\bar{c}\in C_{\bar{\tau}_{0}}$ and all $j\in\{1,\ldots,k\},\bar{c}_{j}\leq f({0},{t})$ or $\bar{c}_{j}=\infty$. Notice that one can choose for $\bar{\sigma}_{0}$ a set of witnesses derived from those of $\sigma_{0}$ having $h$ as prefix: $\textsf{Wit}_{\bar{\sigma}_{0}}=\{\bar{\pi}\in\textsf{Play}_{\bar{G}}\mid\bar{h}\bar{\pi}\in\textsf{Wit}_{\sigma_{0}}\}$. In particular, there exists $\bar{\rho}\in\textsf{Wit}_{\bar{\sigma}_{0}}$ such that $\rho=\bar{h}\bar{\rho}$. By (8) and (9), we have that $\textsf{cost}_{T_{i}}(\bar{\rho})>f({0},{t})$ ($\textsf{cost}_{T_{i}}(\bar{\rho})$ and $\textsf{cost}_{T_{i}}(\rho)$ differ by $w(h)$). As $\bar{c}_{i}\leq f({0},{t})$, it follows that $\bar{\tau}_{0}\prec\bar{\sigma}_{0}$.
We now want to transfer the previous solution $\bar{\tau}_{0}$ in $\bar{G}$ to the subgame $G_{|h}$ in a way to apply Lemma 3.1 and thus obtain the desired strategy $\sigma^{\prime}_{0}$ in $G$. Recall that $\bar{\sigma}_{0}$ was constructed from $\sigma_{0|h}\in\Sigma_{0|h}$. Let us conversely define the strategy $\tau_{0}\in\Sigma_{0|h}$ from $\bar{\tau}_{0}$: $\tau_{0}(\bar{h}g)=\bar{\tau}_{0}(g)$ for all histories $\bar{h}g\in\textsf{Hist}_{G}$. Moreover, we can choose the following sets of witnesses for $\sigma_{0|h}$ and $\tau_{0}$: $\textsf{Wit}_{\sigma_{0|h}}=\{\bar{h}\bar{\pi}\mid\bar{\pi}\in\textsf{Wit}_{\bar{\sigma}_{0}}\}$ and $\textsf{Wit}_{\tau_{0}}=\{\bar{h}\bar{\pi}\mid\bar{\pi}\in\textsf{Wit}_{\bar{\tau}_{0}}\}$. From $\bar{\tau}_{0}\prec\bar{\sigma}_{0}$, it follows that $\tau_{0}\prec\sigma_{0|h}$ (again, the Pareto-optimal costs of $\bar{\tau}$ and $\tau$ differ by $w(h)$, and so do the ones of $\bar{\sigma}_{0}$ and $\sigma_{0}$). Moreover $\tau_{0}\in\mbox{SPS}({G_{|h}},{B})$ since $h$ visits Player 00’s target. Hence, by Lemma 3.1, the strategy $\sigma^{\prime}_{0}=\sigma_{0}[h\rightarrow\tau_{0}]$ is a solution in $\mbox{SPS}({G},{B})$ and $\sigma_{0}^{\prime}\prec\sigma_{0}$.

### v3

@@ Abstract
In this paper, we deepen the study of two-player Stackelberg games played on graphs in which Player 0 announces a strategy and Player $1$, having several objectives, responds rationally by following plays providing him Pareto-optimal payoffs given the strategy of Player 0. The Stackelberg-Pareto synthesis problem, asking whether Player 0 can announce a strategy which satisfies his objective, whatever the rational response of Player $1$, has been recently investigated for $\omega$-regular objectives. We solve this problem for weighted graph games and quantitative reachability objectives such that Player 0 wants to reach his target set with a total cost less than some given upper bound. We show that it is NEXPTIME-complete, as for Boolean reachability objectives.
@@ 2.2 Stackelberg-Pareto Synthesis Problem
In an SP game, Player 0 wishes to minimize the value of a play with respect to the usual order << on $\rm Nature$} extended to $\overline{{\rm Nature}}$ such that $n<\infty$ for all $n\in{\rm Nature}$. To compare the costs of Player $1$, the following component-wise order is introduced. Let $c,c^{\prime}\in\overline{{\rm Nature}}^{t}$ be two costs, we say that $c\leq c^{\prime}$ if $c_{i}\leq c_{i}^{\prime}$ for all $i\in\{1,\ldots,t\}$. Moreover, we write $c<c^{\prime}$ if $c\leq c^{\prime}$ and $c\neq c^{\prime}$. Notice that the order defined on costs is not total. Given two plays with respective costs $c$ and $c^{\prime}$, if $c<c^{\prime}$, then Player $1$ prefers the play with lower cost $c$.
@@
Figure 1: Arena AA (on the left) – Witness tree (on the right)
@@
Interestingly, the Boolean version of this game does not admit any solution. In this case, given a target, the player’s goal is simply to visit it (and not to minimize the cost to reach it). That is, the Boolean version is equivalent to the quantitative one with all weights and the bound $B$ put to zero. In the example, the play $v_{0}v_{1}v_{2}^{\omega}$ is Pareto-optimal (with visits to $T_{1}$ and $T_{3}$), whatever the strategy of Player 0, and this play does not visit Player 0’s target.
@@ 2.3 Tools
Let $X=\{e\in E\mid w(e)\geq 2\}$ be the set of edges of $A$ with weight at least 2. For each $e\in X$, we replace $e$ by a succession of w(e)1w(e)-1 new vertices linked by $w(e)$ new edges of weight 1, and for each $e\in E\setminus X$, we keep $e$ unmodified. In this way, we get a directed graph $(V^{\prime},E^{\prime})$ with a weight function $w^{\prime}\colon E^{\prime}\rightarrow\{0,1\}$ and with a size $|V^{\prime}|\leq|V|\cdot W$. Notice that if $X=\emptyset$, the arena is already binary. The new arena $A^{\prime}=(V^{\prime},V^{\prime}_{0},V^{\prime}_{1},E^{\prime},v_{0},w^{\prime})$ has the same initial vertex $v_{0}$ as $A$, and a partition $V^{\prime}_{0}\cup V^{\prime}_{1}$ of $V^{\prime}$ such that each new vertex is added [^5] Each new vertex could be added to $V_{1}$ as it has a unique successor. to $V_{0}$, hence $V_{0}\subseteq V^{\prime}_{0}$, $V_{1}=V^{\prime}_{1}$. The new SP game $G^{\prime}=(A^{\prime},T_{0},\ldots,T_{t})$ keeps the same targets as in $G$.
@@ 3.2 Improving a Solution in a Subgame
  • If $h\not\rho$, then $\rho$ is also consistent with $\sigma_{0}^{\prime}$, thus there exists $c^{\prime}\in C_{\sigma_{0}^{\prime}}$ such that $c^{\prime}\leq c=\textsf{cost}({\rho})$ by definition of $C_{\sigma_{0}^{\prime}}$.
@@ 3.3 Eliminating a Cycle in a Witness
Let $\sigma_{0}^{\prime}$ be the strategy defined such that $\sigma^{\prime}_{0}(h)=\sigma_{0}(h)$ if $\rho_{\leq m}\not$ h, and $\sigma^{\prime}_{0}(h)=\sigma_{0}(\rho_{\leq n}h^{\prime})$ if $h=\rho_{\leq m}h^{\prime}$. Then $\sigma^{\prime}_{0}$ is a solution in $\mbox{SPS}({G},{B})$ such that $\sigma_{0}^{\prime}\preceq\sigma_{0}$. Moreover, if $\sigma_{0}^{\prime}\simeq\sigma_{0}$, then $\sigma^{\prime}_{0}$ has a witness set $\textsf{Wit}_{\sigma^{\prime}_{0}}$ such that $\textsf{length}({\textsf{Wit}_{\sigma^{\prime}_{0}}})<\textsf{length}({\textsf{Wit}_{\sigma_{0}}})$.
Given a witness $\rho$, the first condition of the lemma means that $\rho_{[m,n]}$ is a cycle and that it appears before the last visit of a target by $\rho$. The second one says that $\rho_{\leq m}\sim\rho_{\leq n}$, i.e., no new target is visited and no branching point is crossed from history $\rho_{\leq m}$ to history $\rho_{\leq n}$. The third one says that if $\rho_{\leq m}$ does not visit Player 0’s target, then the cycle $\rho_{[m,n]}$ must have a null weight. The new strategy $\sigma^{\prime}_{0}$ is obtained from $\sigma_{0}$ by playing after $\rho_{\leq m}$ as playing after $\rho_{\leq n}$ (thus deleting the cycle $\rho_{[m,n]}$). See Figure 3.
[…]
@@
We first prove that $\sigma^{\prime}_{0}\preceq\sigma_{0}$. Let $c\in C_{\sigma_{0}}$ and $\pi\in\textsf{Wit}_{\sigma_{0}}$ be a witness with $\textsf{cost}({\pi})=c$. Let us prove that there exists $c^{\prime}\in C_{\sigma^{\prime}_{0}}$ such that $c^{\prime}\leq c$. If $g\not\pi$, then $\pi\in\textsf{Play}_{\sigma^{\prime}_{0}}$ by definition of $\sigma^{\prime}_{0}$, and thus there exists $c^{\prime}\in C_{\sigma^{\prime}_{0}}$ such that $c^{\prime}\leq c$ by definition of $C_{\sigma^{\prime}_{0}}$. If $g\pi$, as $g\sim h$ by hypothesis, we have that $\textsf{Wit}_{\sigma_{0}}(g)=\textsf{Wit}_{\sigma_{0}}(h)$ and no new target is visited from $g$ to $h$. It follows that $h\pi$ and $\bar{\pi}\in\textsf{Play}_{\sigma^{\prime}_{0}}$ with $\textsf{cost}({\bar{\pi}})=\textsf{cost}({\pi})-w(\rho_{[m,n]})\leq\textsf{cost}({\pi})$. By definition of $C_{\sigma^{\prime}_{0}}$, there exists $c^{\prime}\in C_{\sigma^{\prime}_{0}}$ such that $c^{\prime}\leq\textsf{cost}({\bar{\pi}})$, and therefore $c^{\prime}\leq\textsf{cost}({\pi})=c$. Hence $\sigma^{\prime}_{0}\preceq\sigma_{0}$.
@@
  • In the second case, we consider $\pi\in\textsf{Play}_{\sigma_{0}}$ such that $\bar{\pi}=\pi^{\prime}$. Notice that $\textsf{cost}({\pi})=c^{\prime}$ because $\rho_{[m,n]}$ has a null weight and $g\sim h$ (in particular, $\textsf{cost}({g})=\textsf{cost}({h})$, that is, no new Player 1’s target is visited from $g$ to $h$). We get that $c^{\prime}=\textsf{cost}({\pi})\in C_{\sigma_{0}}$ since $\sigma^{\prime}_{0}\preceq\sigma_{0}$. It follows that $\textsf{val}(\pi)\leq B$ as $\sigma_{0}\in\mbox{SPS}({G},{B})$. As $g\sim h$ (in particular, $\textsf{val}(g)=\textsf{val}(h)$), as $\textsf{val}(g)=\infty$, $\pi$ visits Player 0’s target after $h$, thus not along $\rho_{[m,n]}$. It follows that $\textsf{val}(\pi^{\prime})\leq B$. Therefore $\sigma^{\prime}_{0}\in\mbox{SPS}({G},{B})$.
@@ 4 Bounding Pareto-Optimal Payoffs
The theorem is proved by induction on the dimension $t$, with the calculation of a function f(B,t)f({B},{t}) depending on both $B$ and $t$, that bounds the components $c^{\prime}_{i}\neq\infty$. That is, in Theorem 4.1, Equation (2) is replaced by
@@ 4.1 Dimension One
Notice that f(B,1)f({B},{1}) satisfies (4) when $t=1$. Indeed, we have
@@ 4.2 Bounding the Pareto-Optimal Costs by Induction
The idea of the proof is as follows. If there exists $c\in C_{\sigma_{0}}$ such that for some $i\in\{1,\ldots,t+1\}$, $c_{i}$ does not satisfy (7), then we consider a witness $\rho$ with $\textsf{cost}({\rho})=c$ and the history $h$ of minimal length such that $h\rho$ and $w(h)=\max\{c_{min},B\}+1$. It follows that the subgame $G_{|h}$ has a smaller dimension and we can thus apply the induction hypothesis in $G_{|h}$ with $B=0$ as $h$ has already visited Player 0’s target. Hence, by (3), we get a better solution in the subgame $G_{|h}$, and then a better solution in the whole game $G$ by Lemma 3.1.
@@
Notice that $h$ exists by definition of $c_{i}$ and as the arena is binary. As $\sigma_{0}\in\mbox{SPS}(G,B)$, by definition of $h$, Player 0’s target and at least one target of Player $1$ are visited by $\bar{h}$ such that $\Lasthh=\bar{h}\Last{h}$. Therefore the subgame $G_{|h}$ has dimension $k\leq t$. Notice that $k>0$ by definition of $c_{i}$, see (8).
@@
  • and with the targets visited by $\bar{h}$ removed from Player $1$’s set of targets, and if $k<t$ with tkt-k additional targets that are copies of some remaining (non-empty) targets in a way to have exactly $t$ targets for Player $1$.
@@
We can thus apply the induction hypothesis: by Equation (3), there exists $\bar{\tau}_{0}\in\mbox{SPS}({\bar{G}},{0})$ without cycles such that $\bar{\tau}_{0}\preceq\bar{\sigma}_{0}$ and for all $\bar{c}\in C_{\bar{\tau}_{0}}$ and all $j\in\{1,\ldots,k\},\bar{c}_{j}\leq f({0},{t})$ or $\bar{c}_{j}=\infty$.
Notice that one can choose for $\bar{\sigma}_{0}$ a set of witnesses derived from those of $\sigma_{0}$ having $h$ as prefix: $\textsf{Wit}_{\bar{\sigma}_{0}}=\{\bar{\pi}\in\textsf{Play}_{\bar{G}}\mid\bar{h}\bar{\pi}\in\textsf{Wit}_{\sigma_{0}}\}$. In particular, there exists $\bar{\rho}\in\textsf{Wit}_{\bar{\sigma}_{0}}$ such that $\rho=\bar{h}\bar{\rho}$. By (8) and (9), we have that $\textsf{cost}_{T_{i}}(\bar{\rho})>f({0},{t})$ ($\textsf{cost}_{T_{i}}(\bar{\rho})$ and $\textsf{cost}_{T_{i}}(\rho)$ differ by $w(h)$). As $\bar{c}_{i}\leq f({0},{t})$, it follows that $\bar{\tau}_{0}\prec\bar{\sigma}_{0}$.
We now want to transfer the previous solution $\bar{\tau}_{0}$ in $\bar{G}$ to the subgame $G_{|h}$ in a way to apply Lemma 3.1 and thus obtain the desired strategy $\sigma^{\prime}_{0}$ in $G$. Recall that $\bar{\sigma}_{0}$ was constructed from $\sigma_{0|h}\in\Sigma_{0|h}$. Let us conversely define the strategy $\tau_{0}\in\Sigma_{0|h}$ from $\bar{\tau}_{0}$: $\tau_{0}(\bar{h}g)=\bar{\tau}_{0}(g)$ for all histories $\bar{h}g\in\textsf{Hist}_{G}$. Moreover, we can choose the following sets of witnesses for $\sigma_{0|h}$ and $\tau_{0}$: $\textsf{Wit}_{\sigma_{0|h}}=\{\bar{h}\bar{\pi}\mid\bar{\pi}\in\textsf{Wit}_{\bar{\sigma}_{0}}\}$ and $\textsf{Wit}_{\tau_{0}}=\{\bar{h}\bar{\pi}\mid\bar{\pi}\in\textsf{Wit}_{\bar{\tau}_{0}}\}$. From $\bar{\tau}_{0}\prec\bar{\sigma}_{0}$, it follows that $\tau_{0}\prec\sigma_{0|h}$ (again, the Pareto-optimal costs of $\bar{\tau}$ and $\tau$ differ by $w(h)$, and so do the ones of $\bar{\sigma}_{0}$ and $\sigma_{0}$). Moreover $\tau_{0}\in\mbox{SPS}({G_{|h}},{B})$ since $h$ visits Player 0’s target. Hence, by Lemma 3.1, the strategy $\sigma^{\prime}_{0}=\sigma_{0}[h\rightarrow\tau_{0}]$ is a solution in $\mbox{SPS}({G},{B})$ and $\sigma_{0}^{\prime}\prec\sigma_{0}$.